OpenStudy (anonymous):

Check my answers? GIving medal: 7. What is the slope of the line that passes throught the pair of points (1, 7) and (10, 1)? a. 3/2 b. -2/3 c. -3/2 d. 2/3 <-- 8. What is the slope of the line that passes through the points (-5.5, 6.1) and (-2.5, 3.1)? a. -1 b. 1 c. -3 <-- d. d 9. what is the slope of the line that passes through the pair of points (-7/3, -3) and (-5, 5/2)? a. 6/22 <-- b. -6/22 c. 22/6 d. -22/6

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red} 1}\quad ,&{\color{blue} 7})\quad &({\color{red} {10}}\quad ,&{\color{blue} 1})\\ &({\color{red} {-5.5}}\quad ,&{\color{blue} {6.1}})\quad &({\color{red} {-2.5}}\quad ,&{\color{blue} {3.1}})\\ &({\color{red}{ -\frac{7}{3}}}\quad ,&{\color{blue} {-3}})\quad &({\color{red} {-5}}\quad ,&{\color{blue}{ \frac{5}{2}}}) \end{array} \\\quad \\ slope = {\color{green} m}= \cfrac{rise}{run} \implies \cfrac{{\color{blue} y_2}-{\color{blue} y_1}}{{\color{red} x_2}-{\color{red} x_1}}\)

OpenStudy (anonymous):

oh okay! thanks!

OpenStudy (anonymous):

what one did he do for you

OpenStudy (anonymous):

7 - b 8 - a 9 - -33\16